OpenStudy (anonymous):

A tennis player hits a 58.0 {\rm g} tennis ball so that it goes straight up and reaches a maximum height of 7.35 {\rm m} .How much work does gravity do on the ball on the way up?

OpenStudy (anonymous):

What is the definition of work?

OpenStudy (anonymous):

Work is defined as\[W = F \cdot d\]The force of gravity is expressed as\[F_g = m \cdot g\]Therefore, the work done by gravity is\[W = m \cdot g \cdot d = 58 \cdot 9.81 \cdot 7.35 \]